Radiation intensity of a dipole depends strongly on frequency. If at a frequency f, the intensity of radiation is 'I'. Then at a frequency of f/2, the intensity will be


1
Expert's answer
2021-01-11T11:59:53-0500

We know that radiation intensity of dipole is directly proportional to the power four of frequency i.e


"I=kf^4"

Where I is intensity and f is frequency.If frequency become f/2 ,then intensity become


"I'=k\bigg(\frac{f}{2}\bigg)^4\\I'=k\bigg(\frac{f^4}{16}\bigg)\\I'=\bigg(\frac{I}{16}\bigg)"

This Radiation intensity of dipole become I/16.